OpenStudy (anonymous):

Solve for x: |3x + 2| = 2

OpenStudy (jamesj):

So, for the record newmonei..., what is the definition of |x|?

OpenStudy (anonymous):

3x + 2 = 2 or 3x + 2 = -2 so : 3x = 0 x= 0 or 3x = - 4 x = -4/3 so x = 0 or x = -4/3